### v1

@@ 4.1 Computing the Projection Step
1:Compute $\hat{\mathbf{A}}={\mathbf{A}}_{S}{\mathbf{\Pi}}^{\scriptscriptstyle{\top}}$, where ${\mathbf{\Pi}}\in{\mathds{R}}^{\tau\times s}$ is a random sketching matrix;
@@
For ${\mathbf{A}}_{S}\in{\mathds{R}}^{s\times n}$ and $\lambda\geq 0$, if ${\mathbf{\Pi}}=\sqrt{\frac{n}{\tau}}{\mathbf{I}}_{T}{\mathbf{Q}}$ where ${\mathbf{Q}}$ is the RHT and $T$ is a uniformly random set of size $\tau\geq C(d_{\lambda}({\mathbf{A}}_{S})+\log(n/\delta))\log(d_{\lambda}({\mathbf{A}}_{S})/\delta)$, then with probability $1-\delta$ we have $\kappa({\mathbf{R}}^{-{\scriptscriptstyle{\top}}}\big{[}{\mathbf{A}}_{S}\ \sqrt{\lambda}{\mathbf{I}}\big{]})\leq 2$, and after $O(\log(\kappa_{S}/\epsilon))$ iterations of LSQR on (4.1), we get $\tilde{\mathbf{w}}$ such that $\|\tilde{\mathbf{w}}-{\mathbf{w}}_{t}\|\leq\epsilon\|{\mathbf{w}}_{t}\|$, where $\kappa_{S}=\max\{1,\sqrt{\lambda}/\sigma_{\min}({\mathbf{A}}_{S})\}$.
@@
Since ${\mathbf{w}}_{t}={\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+\lambda{\mathbf{I}})^{-1}{\mathbf{r}}_{t}$ and ${\mathbf{v}}_{t}=\sqrt{\lambda}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+\lambda{\mathbf{I}})^{-1}{\mathbf{r}}_{t}$, it follows that ${\mathbf{w}}_{t}=\frac{1}{\sqrt{\lambda}}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}{\mathbf{v}}_{t}$. This implies that $\|{\mathbf{v}}_{t}\|\leq\kappa_{S}\|{\mathbf{w}}_{t}\|$, and so $\|\tilde{\mathbf{w}}-{\mathbf{w}}_{t}\|\leq 2\epsilon(1+\kappa_{S})\|{\mathbf{w}}_{t}\|$. Adjusting $\epsilon$ appropriately concludes the proof. ∎
Constructing the preconditioner ${\mathbf{R}}$ takes $O(T_{\mathrm{sketch}}+\tau s^{2}+s^{3})$ operations, where $T_{\mathrm{sketch}}$ represents the cost of the matrix product ${\mathbf{A}}_{S}{\mathbf{\Pi}}^{\scriptscriptstyle{\top}}$. For example, when ${\mathbf{\Pi}}$ is the SRHT, as in Lemma 4.1, then $T_{\mathrm{sketch}}=O(ns\log n)$. Thus, setting $\tau=O(s\log s)$, the overall cost of solving the projection step to within $\epsilon$ relative accuracy takes no more than $O(ns\log n+s^{3}\log s)$ operations for constructing ${\mathbf{R}}$, followed by $O(ns\log(\kappa_{S}/\epsilon))$ operations for running LSQR.
@@ 4.2 Block Memoization
Consider matrix ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$, parameters $\bar{\lambda}\geq\lambda>0$, and a probability distribution ${\mathcal{D}}$ over subsets of $[m]$, such that the regularized projection matrix ${\mathbf{P}}_{{\lambda},S}\coloneqq{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}{\mathbf{A}}_{S}$ satisfies:
@@
for some $c>0$. Let =$\mathcal{B}=\{S_{i}\}_{i=1}^{B}$ be a collection of $B$ independent samples from $\mathcal{D}$. If $B\geq\frac{4(7c+13)\bar{\lambda}}{3c\lambda}\log(2n/\delta)$, then with probability $1-\delta$, the collection ${\mathcal{B}}$ satisfies:
@@
Lemma 4.3 (Matrix Bernstein).
Let ${\mathbf{Z}}_{1},\ldots,{\mathbf{Z}}_{B}$ be independent random symmetric $n\times n$ matrices such that $\frac{1}{B}\sum_{j}{\mathds{E}}[{\mathbf{Z}}_{j}]=\bar{{\mathbf{Z}}}$ and $\|\frac{1}{B}\sum_{j}{\mathds{E}}[({\mathbf{Z}}_{j}-\bar{{\mathbf{Z}}})^{2}]\|\leq\sigma^{2}$. Suppose we have $\|{\mathbf{Z}}_{j}-{\mathds{E}}[{\mathbf{Z}}_{j}]\|\leq R$ holds for all $j\in[B]$. Then for any $\epsilon\geq 0$,
Proof of Theorem 4.2.
Let ${\mathbf{Z}}_{j}\coloneqq\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}{\mathbf{P}}_{{\lambda},S_{j}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}$ be a normalized form of the regularized projection matrix ${\mathbf{P}}_{{\lambda},S_{j}}$ associated with subset $S_{j}$ from $\mathcal{B}$. Then, we have ${\mathds{E}}_{S_{j}\sim\mathcal{D}}[{\mathbf{Z}}_{j}]=\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\bar{{\mathbf{P}}}_{{\lambda}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}$. In order to apply matrix Bernstein and give a concentration result, we will bound $\|{\mathbf{Z}}_{j}-\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\bar{{\mathbf{P}}}_{{\lambda}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\|$ and $\|{\mathds{E}}_{S_{j}\sim\mathcal{D}}[({\mathbf{Z}}_{j}-\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\bar{{\mathbf{P}}}_{{\lambda}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2})^{2}]\|$. For the second term, we have
$\|{\mathds{E}}_{S_{j}\sim\mathcal{D}}[({\mathbf{Z}}_{j}-\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\bar{{\mathbf{P}}}_{{\lambda}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2})^{2}]\|=$ $~{}\|\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}{\mathds{E}}[({\mathbf{P}}_{{\lambda},S_{j}}-\bar{{\mathbf{P}}}_{{\lambda}})\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger}({\mathbf{P}}_{{\lambda},S_{j}}-\bar{{\mathbf{P}}}_{{\lambda}})]\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\|$
$=$ $~{}\|\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}({\mathds{E}}[{\mathbf{P}}_{{\lambda},S_{j}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger}{\mathbf{P}}_{{\lambda},S_{j}}]-\bar{{\mathbf{P}}}_{{\lambda}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger}\bar{{\mathbf{P}}}_{{\lambda}})\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\|$
$=$ $~{}\|{\mathds{E}}[\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}{\mathbf{P}}_{{\lambda},S_{j}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger}{\mathbf{P}}_{{\lambda},S_{j}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}]-\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\bar{{\mathbf{P}}}_{{\lambda}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\|$
$\leq$ $~{}\|{\mathds{E}}[(\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}{\mathbf{P}}_{{\lambda},S_{j}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2})^{2}]\|+1=\nu+1,$
where the last step follows from the definition of $\nu=\nu({\mathbf{A}},{\mathcal{D}},\lambda)$ in (2.1). According to Theorem 3.7, under the same assumption we have $\nu\leq\frac{2\bar{\lambda}}{c{\lambda}}$. By applying this result we obtain the bound $\|{\mathds{E}}_{S_{j}\sim\mathcal{D}}[({\mathbf{Z}}_{j}-\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\bar{{\mathbf{P}}}_{{\lambda}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2})^{2}]\|\leq\frac{2\bar{\lambda}}{c{\lambda}}+1=:\sigma^{2}$.
Next, to obtain the $R$ term in Bernstein’s inequality, notice that according to our assumption, $\bar{\mathbf{P}}_{{\lambda}}^{\dagger}\preceq\frac{1}{c}({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}}+\bar{\lambda}{\mathbf{I}})({\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}})^{\dagger}=:{\mathbf{\Sigma}}$, thus we have
$\|{\mathbf{Z}}_{j}\|\leq\|{\mathbf{\Sigma}}^{1/2}{\mathbf{P}}_{{\lambda},S_{j}}{\mathbf{\Sigma}}^{1/2}\|=\|{\mathbf{\Sigma}}^{1/2}{\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{I}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}{\mathbf{I}}_{S}{\mathbf{A}}{\mathbf{\Sigma}}^{1/2}\|.$
Letting ${\mathbf{A}}={\mathbf{U}}{\mathbf{D}}{\mathbf{V}}^{\scriptscriptstyle{\top}}$ be the compact SVD of ${\mathbf{A}}$, note that ${\mathbf{\Sigma}}{\mathbf{A}}^{\scriptscriptstyle{\top}}=\frac{1}{c}{\mathbf{V}}({\mathbf{D}}^{2}+\bar{\lambda}{\mathbf{I}})^{1/2}{\mathbf{U}}^{\scriptscriptstyle{\top}}$. So, additionally observing that ${\mathbf{U}}({\mathbf{D}}^{2}+\bar{\lambda}{\mathbf{I}}){\mathbf{U}}^{\scriptscriptstyle{\top}}\preceq{\mathbf{A}}{\mathbf{A}}^{\scriptscriptstyle{\top}}+\bar{\lambda}{\mathbf{I}}\preceq\frac{\bar{\lambda}}{\lambda}({\mathbf{A}}{\mathbf{A}}^{\scriptscriptstyle{\top}}+\lambda{\mathbf{I}})$, we get
$\|{\mathbf{Z}}_{j}\|$ $\leq\frac{1}{c}\|{\mathbf{V}}({\mathbf{D}}^{2}+\bar{\lambda}{\mathbf{I}})^{1/2}{\mathbf{U}}^{\scriptscriptstyle{\top}}{\mathbf{I}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1}{\mathbf{I}}_{S}{\mathbf{U}}({\mathbf{D}}^{2}+\bar{\lambda}{\mathbf{I}})^{1/2}{\mathbf{V}}^{\scriptscriptstyle{\top}}\|$
$=\frac{1}{c}\|({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1/2}{\mathbf{I}}_{S}{\mathbf{U}}({\mathbf{D}}^{2}+\bar{\lambda}{\mathbf{I}}){\mathbf{U}}^{\scriptscriptstyle{\top}}{\mathbf{I}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1/2}\|$
$\leq\frac{\bar{\lambda}}{c\lambda}\|({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1/2}{\mathbf{I}}_{S}({\mathbf{A}}{\mathbf{A}}^{\scriptscriptstyle{\top}}+\lambda{\mathbf{I}}){\mathbf{I}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+{\lambda}{\mathbf{I}})^{-1/2}\|=\frac{\bar{\lambda}}{c\lambda}.$
This gives us the $R$ term in Bernstein’s inequality:
$\|{\mathbf{Z}}_{j}-\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\bar{{\mathbf{P}}}_{{\lambda}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\|\leq\|{\mathbf{Z}}_{j}\|+\|\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\bar{{\mathbf{P}}}_{{\lambda}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\|\leq\frac{\bar{\lambda}}{c{\lambda}}+1\leq\frac{(c+1)\bar{\lambda}}{c{\lambda}}=:R.$
Applying Lemma 4.3 to matrices $\{{\mathbf{Z}}_{j}\}_{j=1}^{B}$ with parameters $R=\frac{(c+1)\bar{\lambda}}{c{\lambda}}$ and $\sigma^{2}=\frac{2\bar{\lambda}}{c{\lambda}}+1$, we conclude that if we set $B\geq\frac{4(7c+13)\bar{\lambda}}{3c{\lambda}}\log(2n/\delta)$ and $\bar{\mathbf{Z}}=\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}\bar{{\mathbf{P}}}_{{\lambda}}\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger/2}$, then
$\Pr\left(\big{\|}\frac{1}{B}\sum_{j=1}^{B}{\mathbf{Z}}_{j}-\bar{\mathbf{Z}}\big{\|}\geq\frac{1}{2}\right)\leq 2n\cdot\exp\left(-\frac{B/8}{\frac{2\bar{\lambda}}{c{\lambda}}+1+\frac{(c+1)\bar{\lambda}}{6c{\lambda}}}\right)\leq\delta.$
So, the average $\hat{\mathbf{Z}}:=\frac{1}{B}\sum_{j=1}^{B}{\mathbf{Z}}_{j}$ satisfies $\hat{\mathbf{Z}}\succeq\bar{\mathbf{Z}}-\frac{1}{2}{\mathbf{I}}$ with probability $1-\delta$. Note that $\bar{\mathbf{Z}}$ is a projection onto a subspace that contains the range of $\hat{\mathbf{Z}}$, so applying $\bar{\mathbf{Z}}$ on both sides of that inequality we get $\hat{\mathbf{Z}}=\bar{\mathbf{Z}}\hat{\mathbf{Z}}\bar{\mathbf{Z}}\succeq\bar{\mathbf{Z}}\bar{\mathbf{Z}}\bar{\mathbf{Z}}-\frac{1}{2}\bar{\mathbf{Z}}\bar{\mathbf{Z}}=\frac{1}{2}\bar{\mathbf{Z}}$. This gives
$\frac{1}{B}\sum_{j=1}^{B}{\mathbf{P}}_{{\lambda},S_{j}}\succeq\bar{{\mathbf{P}}}_{{\lambda}}^{1/2}\hat{\mathbf{Z}}\bar{{\mathbf{P}}}_{{\lambda}}^{1/2}\succeq\frac{1}{2}\bar{{\mathbf{P}}}_{{\lambda}}^{1/2}\bar{\mathbf{Z}}\bar{{\mathbf{P}}}_{{\lambda}}^{1/2}=\frac{1}{2}\bar{{\mathbf{P}}}_{{\lambda}}.$
Combining Theorem 4.2 with the analysis of $\mu$ and $\nu$ in Theorems 3.1 and 3.7, as well as the acceleration analysis from Theorem 2.1, we can now recover the convergence guarantee for Kaczmarz++ (Algorithm 1).
@@
Suppose ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$ is transformed by RHT, i.e., $\bar{\mathbf{A}}={\mathbf{Q}}{\mathbf{A}}$. Let $\sigma_{1}\geq\sigma_{2}\geq...$ be ${\mathbf{A}}$’s singular values. Given $\delta\in(0,1)$ and $C\log(m/\delta)\leq k<\mathrm{rank}({\mathbf{A}})$, let $\lambda=\frac{1}{m}\sum_{i>k}\sigma_{i}^{2}$. If ${\mathcal{B}}$ consists of $O(\frac{m}{k}\log(n/\delta))$ uniformly random sets from (${[m]\choose s}$ for $s\geq Ck\log(m\bar{\kappa}_{k})$ and ${\mathcal{U}}({\mathcal{B}})$ is the uniform distribution over ${\mathcal{B}}$, then with prob. $1-\delta$:
$\mu\big{(}\bar{\mathbf{A}},{\mathcal{U}}({\mathcal{B}}),\lambda\big{)}\geq\frac{k}{4m\bar{\kappa}_{k}^{2}},\quad\nu\big{(}\bar{\mathbf{A}},{\mathcal{U}}({\mathcal{B}}),\lambda\big{)}\leq\frac{8m}{k}\quad\text{and}\quad\bar{\rho}\big{(}\bar{\mathbf{A}},{\mathcal{U}}({\mathcal{B}}),\lambda\big{)}\geq\frac{k}{6m\bar{\kappa}_{k}}.$
Thus, via Theorem 2.1, Kaczmarz++ with $\rho=\bar{\rho}/2$ and $\eta=\frac{1}{2\nu}$ satisfies:
${\mathds{E}}\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|^{2}\leq 8\Big{(}1-\frac{k}{24m\bar{\kappa}_{k}}\Big{)}^{t}\|{\mathbf{x}}_{0}-{\mathbf{x}}^{*}\|^{2}.$

### v2

@@ 4.1 Computing the Projection Step
1:Compute $\hat{\mathbf{A}}={\mathbf{A}}_{S}{\mathbf{\Pi}}^{\scriptscriptstyle{\top}}$, where ${\mathbf{\Pi}}\in{\mathds{R}}^{\tau\times n}$ is a random sketching matrix;
@@
For ${\mathbf{A}}_{S}\in{\mathds{R}}^{s\times n}$ and $\lambda\geq 0$, if ${\mathbf{\Pi}}=\sqrt{\frac{n}{\tau}}{\mathbf{I}}_{T}{\mathbf{Q}}$ where ${\mathbf{Q}}$ is the RHT and $T$ is a uniformly random set of size $\tau\geq C(d_{\lambda}({\mathbf{A}}_{S})+\log(n/\delta))\log(d_{\lambda}({\mathbf{A}}_{S})/\delta)$, then with probability $1-\delta$ we have $\kappa({\mathbf{R}}^{-{\scriptscriptstyle{\top}}}\big{[}{\mathbf{A}}_{S}\ \sqrt{\lambda}{\mathbf{I}}\big{]})\leq 2$, and after $O(\log(1/\epsilon))$ iterations of LSQR on (4.1), we get $\tilde{\mathbf{w}}$ such that $\|\tilde{\mathbf{w}}-{\mathbf{w}}_{t}\|\leq\epsilon\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|$.
@@
Since ${\mathbf{w}}_{t}={\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+\lambda{\mathbf{I}})^{-1}{\mathbf{r}}_{t}$, ${\mathbf{v}}_{t}=\sqrt{\lambda}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+\lambda{\mathbf{I}})^{-1}{\mathbf{r}}_{t}$, and ${\mathbf{r}}_{t}={\mathbf{A}}_{S}({\mathbf{x}}_{t}-{\mathbf{x}}^{*})$, we can bound $\|{\mathbf{w}}_{t}\|^{2}$ and $\|{\mathbf{v}}_{t}\|^{2}$ separately as
$\|{\mathbf{w}}_{t}\|^{2}$ $=\|{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+\lambda{\mathbf{I}})^{-1}{\mathbf{A}}_{S}({\mathbf{x}}_{t}-{\mathbf{x}}^{*})\|^{2}\leq\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|^{2}$
and
$\|{\mathbf{v}}_{t}\|^{2}$ $=\lambda({\mathbf{x}}_{t}-{\mathbf{x}}^{*})^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+\lambda{\mathbf{I}})^{-2}{\mathbf{A}}_{S}({\mathbf{x}}_{t}-{\mathbf{x}}^{*})$
$\leq({\mathbf{x}}_{t}-{\mathbf{x}}^{*})^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+\lambda{\mathbf{I}})^{-1}{\mathbf{A}}_{S}({\mathbf{x}}_{t}-{\mathbf{x}}^{*})\leq\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|^{2}.$
Thus $\|\tilde{\mathbf{w}}-{\mathbf{w}}_{t}\|\leq 2\sqrt{2}\epsilon\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|$. Adjusting $\epsilon$ appropriately concludes the proof. ∎
Constructing the preconditioner ${\mathbf{R}}$ takes $O(T_{\mathrm{sketch}}+\tau s^{2}+s^{3})$ operations, where $T_{\mathrm{sketch}}$ represents the cost of the matrix product ${\mathbf{A}}_{S}{\mathbf{\Pi}}^{\scriptscriptstyle{\top}}$. For example, when ${\mathbf{\Pi}}$ is the SRHT, as in Lemma 4.1, then $T_{\mathrm{sketch}}=O(ns\log n)$. Thus, setting $\tau=O(s\log s)$, the overall cost of solving the projection step to within $\epsilon$ relative accuracy takes no more than $O(ns\log n+s^{3}\log s)$ operations for constructing ${\mathbf{R}}$, followed by $O(ns\log(1/\epsilon))$ operations for running LSQR.
@@ 4.2 Block Memoization
[…]
@@
for some $c\in(0,1]$. Assume that ${\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}{\mathbf{A}}_{S}\preceq\alpha{\mathbf{A}}^{\scriptscriptstyle{\top}}{\mathbf{A}}$ holds with probability $1-\delta^{\prime}$ for some $\alpha\in[0,1]$ and $\delta^{\prime}\in[0,\alpha/\|\bar{{\mathbf{P}}}_{\lambda}^{\dagger}\|]$. Let =$\mathcal{B}=\{S_{i}\}_{i=1}^{B}$ be a collection of $B$ independent samples from $\mathcal{D}$. If $B\geq\frac{40}{c}(1+\alpha\frac{\bar{\lambda}}{\lambda})\cdot\log(2n/\delta)$ for some $\delta\in(0,1)$, then with probability $1-B\delta^{\prime}-\delta$, the collection ${\mathcal{B}}$ satisfies:
@@
The proof of Theorem 4.2 follows along similar lines as the proof of Theorem 3.8, since it primarily requires bounding the variance of each term ${\mathbf{P}}_{{\lambda},S_{j}}$, so that we can apply a matrix concentration inequality (details in Supplement D).
The following corollary combines Theorem 4.2 with Theorem 3.1, while also incorporating Lemma 3.9 to account for over-determined systems.
@@
[…]
By further choosing $\lambda=\frac{k}{m}\bar{\lambda}$, the required number of blocks is:
$B\geq 80\left(\frac{\min\{m,9r\log(mn\bar{\kappa}_{k})\}}{k}+1\right)\cdot\log(2n/\delta).$